OpenStudy (anonymous):

Evaluate the trigonometric limit. lim x→0 sin(πx)^2/7x^2

OpenStudy (anonymous):

\[\lim_{x \rightarrow 0} \sin^2(\pi x)/7x^2 \]

OpenStudy (anonymous):

please do not use l'hopital

OpenStudy (anonymous):

Use L'Hopital's Rule.

OpenStudy (anonymous):

Ah, why not?

OpenStudy (anonymous):

can't we haven't learned that

OpenStudy (anonymous):

that is not how the teacher wants it solved

OpenStudy (anonymous):

What lesson are you learning? It seems to me that L'Hopital is the only way to solve this.

OpenStudy (anonymous):

one second

OpenStudy (anonymous):

One second?

OpenStudy (anonymous):

looking at my notes

OpenStudy (anonymous):

we have been learning implicits but i don't see how it applies

OpenStudy (anonymous):

and just general trig differentiation

OpenStudy (anonymous):

Trig identies could be used.

OpenStudy (anonymous):

Then multiply by conjugate.

OpenStudy (anonymous):

Should work out smoothly...

OpenStudy (anonymous):

@Dido525 What about denominator?

OpenStudy (anonymous):

Hmm let me try...

OpenStudy (kropot72):

For very small angles sin y = y \[\lim_{x \rightarrow 0}\sin (\pi x)^{2}=(\pi x)^{2}\] \[\lim_{x \rightarrow 0}\frac{(\pi x)^{2}}{7x ^{2}}=?\]

OpenStudy (kropot72):

\[\frac{(\pi x)^{2}}{7x ^{2}}=\frac{\ \pi ^{2}}{7}\]

OpenStudy (anonymous):

i plugged in wolfram and it said\[ \pi^2/7 \]

OpenStudy (raden):